OpenStudy (anonymous):

A 1-kg bomb initially at rest explodes into 1-kg and 2-kg fragments. If the kinetic energy of the 1-kg fragment is 600 J, then what is the kinetic energy of the 2-kg fragment?

OpenStudy (anonymous):

From the conservation of linear momentum: Pi=Pf =>0=sqrt(2M1*K.E)+M2*V2=>0=sqrt(2*1*600)+2*V2 =>0=34.64+2*V2 therefore,V2=-34.64/2=-17.32 m/sec hence kinetic energy of 2 Kg mass=1/2M2V2^2=1/2*2*(17.32)^2=299.98J(about)

OpenStudy (anonymous):

Thanks! If you could take a look at my other questions too, I'd really appreciate it

OpenStudy (anonymous):

what's your other question???